OpenStudy (anonymous):

The sum of two numbers is 52. The greater number is 4 more than the smaller number. Which equation can be used to solve for the smaller number?

OpenStudy (anonymous):

Let x be the smaller number. How would you write the larger number in terms of x?

OpenStudy (anonymous):

let larger no. be y then y = x+4(given) and x+y =52 (given) use both to find x

OpenStudy (mathmale):

"The greater number is 4 more than the smaller number" translates into "y=x+4" You now have a "system of linear equations." Solve simultaneously: y=x+4 x+y=52 Methods of solving such a system include (among others): graphing, elimination by addition/subtraction, substition, matrices.